# A practical guide to Digital Micro-mirror Devices (DMDs) for wavefront shaping

### v4

@@ 1 Introduction
Since the advent of adaptive optics, various technologies have been employed to modulate the amplitude and/or phase of light. Early adaptive optics devices, utilized in fields like microscopy and astronomy, offer rapid modulation capable of compensating for the aberrations of optical systems in real-time. However, these devices are constrained by a limited number of actuators, restricting their utility in complex media where a large number of degrees of freedom is essential. Liquid Crystal Spatial Light Modulators (LC-SLMs), which allow for the control of light phase across typically more than a million pixels, have emerged as powerful tools for wavefront shaping in complex media since the seminal work of A. Most and I. Vellekoop in the mid-2000s [1]. Nonetheless, LC-SLMs are hampered by their slow response time, permitting only a modulation speed ranging from a few Hz to a few hundred Hz.
@@
Third-party companies have developed kits tailored for research applications, which include a DMD, a control board, and a software interface. Specifically, Vialux devices [8] offer an FPGA board that enables high-speed modulation by allowing frames to be stored in the device’s memory [9]. However, standard Texas Instruments video projector evaluation modules can also be repurposed into wavefront shaping devices [10], though at a compromised modulation speed. These systems can further be converted into phase or complex field modulators. This is typically achieved by encoding the optical phase into the spatial displacement of binary spatial fringes displayed on the DMD, followed by filtering the high spatial frequencies in the Fourier plane [11]. Such a configuration permits multi-level complex modulation but sacrifices spatial resolution. The implementation and performance of such systems are explored in a separate tutorial [12] and will not be elaborated upon here. For the remainder of this paper, it will be assumed that the DMD is used for complex modulation via such a method.
While other articles exist describing the various aspects of DMDs [13, 14, 10, 15], this tutorial aims to provide a guide for easily setting up a DMD for wavefront shaping applications in complex media. In particular, we provide characterization and validation procedures that require minimal changes compared to typical wavefront shaping setups. More specifically, we place ourselves in a standard experiment where the goal is to shape the complex wavefront impinging on a complex medium and control or measure its output response. This is typically the case for a focusing experiment [1] or for measuring the transmission matrix of a complex medium [16].
@@ 2.1 A 1D model
A significant distinction between liquid crystal modulators and DMDs lies in the geometry of the pixel surface: The pixels of the DMD are not flat, but are mirrors that can be tilted in two different positions around a rotation axis along the diagonal of the square shape of the pixels. This difference gives rise to diffraction effects that can adversely affect both modulation quality and system efficiency. The impact of these diffraction effects is highly dependent on several factors: the wavelength of illumination, the pixel pitch, and both the incident and outgoing angles. Therefore, in conjunction with selecting an appropriate anti-reflection coating, it is crucial to ensure that the pixel pitch is compatible with the specific configuration being used. Texas Instruments offers chips with a variety of pixel pitches $d$, ranging approximately from $5$ to $\sim 25$ µm [18].
To achieve a qualitative understanding of this issue, we consider a 1D array of pixels as illustrated in Fig. 2. Initially, let’s assume that all pixels are in the same state and are illuminated by a plane wave originating from the far field. Under these conditions, the pixelated modulator essentially functions as a grating, with a period $d$ that is equivalent to the pixel pitch. It is important to underscore that these modulators possess a hardware-limited fill factor, typically around 90%. This translates to an effective active pixel size of $d^{\prime}<d$.
In general, a grating gives rise to various diffraction orders with different intensities and angles $\theta_{p}$, as dictated by the grating equation: $\text{sin}(\theta_{p})+\text{sin}(\alpha)=p\lambda/d=p\,\text{sin}(\theta_{D}),$ where $\lambda$ is the wavelength of the light, $\alpha$ is the incident angle, $\theta_{D}=\arcsin\left(\lambda/S\right)$ is the angle of the first diffraction order, and $p$ is an integer value denoting the orders of diffraction. The intensity of the individual diffraction orders is influenced by the response of a single pixel, constituting the unit cell of the grating, and that is governed by $d$, $d^{\prime}$, and its rotation angle in the case of a DMD. Importantly, we can decouple the effects of the grating’s periodicity, which influences the angles of the diffraction orders, from the effects of the response of a single pixel, which shapes the envelope of the angular response.
@@
For a case of normal incidence with an LC-SLM (Liquid Crystal Spatial Light Modulator), we can assume that the response of a single pixel is uniform over its surface. Similarly, in the scenario of a blazed grating, such as a DMD, a linear phase slope is present on each pixel. This is due to the tilt angle $\theta_{B}$ of the mirrors. For an arbitrarily selected incidence angle $\alpha$, a global phase slope is introduced. This results in a trivial shift of the angular diffraction pattern by an angle $\alpha$. The incident angle $\alpha$ has the effect of shifting the entire angular diffraction pattern relative to the case of normal incidence, while the blazed angle $\theta_{B}$ —the tilt angle of the micro-mirrors of the DMD— only shifts the envelope by an angle of $2\theta_{B}$. Whenever the fill factor approaches 100%, i.e. when $d\approx d^{\prime}$, the envelope for a flat grating achieves its maximum at $\theta_{0}=-\alpha$; this corresponds to the angle of the zeroth-order diffraction, and the intensity is comparatively small for all other orders. In these specific conditions, a singular diffraction order is visually perceived, corresponding to the optimum scenario. The addition of a blazed angle $\theta_{B}$ results in both a shift in the envelope and in the position of its maximum $\theta_{0}=2\theta_{B}-\alpha$. In the general case, this position may not align with a diffraction order anymore [13]. We provide in Appendix A. a 1D computation of this effect. A more accurate computation of the far field can be found in  [15].
We represent in Fig. 3 the angular response of a flat grating and a blazed grating for a 1D filling fraction of 95% (correponding to a 2D filling fraction of $\approx 90$%). For a flat grating, the zero-th order contains most of the intensity, the other orders being negligible in comparison. For the blazed grating example shown, we are in a situation close to the worst case scenario: Two diffraction orders have a significant and comparable intensity, and other orders also have non-negligible contributions. In the optimal scenario, where the peak of the envelope corresponds to a diffraction order, it results in a single diffraction order carrying the majority of the energy. This state is achieved when the conditions of the blazed grating equation are fulfilled [19]:
@@
Figure 3: Flat grating vs blazed grating. Far field diffraction patterns for a 1D flat grating (left) and a 1D blazed grating (right) for an input angle of $\alpha=-20^{\circ}$, a filling fraction of 95% (corresponding to a 2D filling fraction of $\approx 90$%), a pixel tilt angle of $\theta_{B}=5^{\circ}$, and a wavelength to pixel pitch ratio $\lambda/d=0.05$. Vertical lines represent the angles of the diffraction orders, peaked at $\theta_{p}$ which satisfy $\sin(\theta_{p})+\sin(\alpha)=p\lambda/d$, and the black dashed curve represents the amplitude of the field which peaks at $\theta_{0}=2\theta_{B}-\alpha$.
@@ 2.2 The 2D case
To analyze more precisely the effect of diffraction in a DMD, one needs to consider the 2D surface of the modulator. We can establish a Cartesian coordinate system on the plane of the DMD, with axes $x$ and $y$ aligned with the pixel sides (refer to Fig. 4a). The pixels are uniformly repeated along these axes. However, a technical challenge arises in that the axis of rotation of the pixels aligns with the pixel diagonals, resulting in a rotation by 45 degrees with respect to the $x$ and $y$ axes. For the convenience of alignment and manipulation of the optical setup, it is preferable to work with the incident and outgoing beams which have the optical axis contained in the horizontal plane, i.e. a plane parallel to the table surface. A straightforward and common solution is to rotate the chip by 45 degrees relative to the horizontal plane, which aligns the pixel axis of rotation to be vertical. This configuration is depicted in Fig. 4b.
Figure 4: 2D grating geometry. (a) Schematic representation of the geometry of the DMD. The incident angle $\alpha$ and the reflection angle $\theta_{0}$ (defined by the peak of the diffraction envelope) are situated within the horizontally plane, illustrated in yellow. (b) Photograph of the DMD chip oriented so that the rotation axis of the pixels is aligned vertically.
with $\vec{k}_{in}$ the incoming wavevector and $\vec{r}$ the position in the modulator plane, represented by the coordinaxte $(x,y)$.
The total phase then reads:
$\phi(x,y)=\phi(x+d,y)=\phi(x,y+d)\,.$ (5)
The new blazed grating equation then reads :
@@
We can quantify how close we are to the ideal case, i.e., when satisfying the blazed this equation, by defining a blazed number $\mu$ as introduced in [20]:
@@
where $\mod{2}$ represents the modulo 2 operation. $\mu$ is maximal and equals $1$ when the blazing equation is satisfied, i.e. when one order of diffraction contains most of the energy, and minimal when we are in the worst-case scenario, i.e. when four orders of diffraction have a significant and equal intensity.
@@
To demonstrate the effect, we first conduct a simulation of a DMD using Python (refer to tutorial and code in [20]) with two pixel pitches of d=7.6d=7.6µm and d=10.8d=10.8µm, under a coherent excitation at $\lambda=633$nm. Fig. 5 shows the estimated blazed number $\mu$ as a function of the angle of incidence $\alpha$ in the horizontal plane, along with the far field diffraction pattern for two distinct incident angles. Fig. 6 shows the blazed number as a function of both the incident angle and the wavelength for pixel pitches of d=7.6d=7.6µm and d=10.8d=10.8µm. It should be noted that the efficiency of diffraction can be altered by adjusting the angle of incidence $\alpha$. However, its impact is relatively confined within an acceptable angular range that aligns with experimental limitations (i.e. for angles far from $\pm 90^{\circ}$). Far-field patterns are centered around the maximum of the envelope for an output angle $2\theta_{B}-\alpha$ (marked by a yellow cross). We see that for small positive values of $\alpha$, the pixel pitch of d=10.8d=10.8µm leads to a blazed number $\mu$ close to $1$. It corresponds in the far field to having one bright order of diffraction close to the maximum of the envelope.
@@
Figure 5: Blazed number and far-field diffraction patterns. Blazing number $\mu$ (Eq.(7)) as a function of the incident angle $\alpha$ (top) for a pixel pitch of d=7.6d=7.6µm (a.) and d=10.8d=10.8µm (b.). Corresponding far-field diffraction patterns (bottom) for two incident angles $\alpha=-15^{\circ}$ and $\alpha=0^{\circ}$. The red cross indicates the maximum of the envelope $\theta_{\text{max}}=2\theta_{B}-\alpha$.
@@
Figure 6: Blazed number as a function of the wavelength and the incident angle. Color map representation of the blazed number for a pixel pitch of d=7.6d=7.6µm (a.) and d=10.8d=10.8µm (b.).
@@
To illustrate the link between the blaze number and the efficiency of the diffraction setup, we compute the ratio of the energy in the brightest diffraction spot to the total diffracted energy for different values of the wavelength, with fixed values for the pixel pitch (7.6 µm) and incident angle (normal incidence). We simulate the DMD with no gap between the pixels (i.e., a filling fraction of 100%) and with 10 pixels in each direction, where all the mirrors of the pixels are in the same state. The results are shown in Fig. 7. We observe that the maxima (resp. minima) of the diffraction efficiency correspond to the maxima (resp. minima) of the blaze number. Note that the actual values of the minimum and maximum diffraction efficiency depend on the parameters of the DMD, such as pixel pitch, filling fraction, resolution, etc.
@@
Figure 7: Blazed number vs diffraction efficiency. Plot of the blazed number and the diffraction efficiency for a normal incidence and a pixel pitch of 7.6 µm as a function of the wavelength. The diffraction efficiency is estimated by comparing the energy in a diffracted limited spot at the location of the brightest diffraction order to the total diffracted energy.

### v5

@@ 1 Introduction
Since the advent of adaptive optics, various technologies have been employed to modulate the amplitude and/or phase of light. Early adaptive optics devices, utilized in fields like microscopy and astronomy, offer rapid modulation capable of compensating for the aberrations of optical systems in real-time. However, these devices are constrained by a limited number of actuators, restricting their utility in complex media where a large number of degrees of freedom is essential. Liquid Crystal Spatial Light Modulators (LC-SLMs), which allow for the control of light phase across typically more than a million pixels, have emerged as powerful tools for wavefront shaping in complex media since the seminal work of A. Mosk and I. Vellekoop in the mid-2000s [1]. Nonetheless, LC-SLMs are hampered by their slow response time, permitting only a modulation speed ranging from a few Hz to a few hundred Hz.
@@
Third-party companies have developed kits tailored for research applications, which include a DMD, a control board, and a software interface. Specifically, Vialux devices [8] offer an FPGA board that enables high-speed modulation by allowing frames to be stored in the device’s memory [9]. However, standard Texas Instruments video projector evaluation modules can also be repurposed into wavefront shaping devices [10], though at a compromised modulation speed. These systems can further be converted into phase or complex field modulators.
While other articles exist describing the various aspects of DMDs [11, 12, 10, 13], this tutorial aims to provide a guide for easily setting up a DMD for wavefront shaping applications in complex media. In particular, we provide characterization and validation procedures that require minimal changes compared to typical wavefront shaping setups. More specifically, we place ourselves in a standard experiment where the goal is to shape the complex wavefront impinging on a complex medium and control or measure its output response. This is typically the case for a focusing experiment [1] or for measuring the transmission matrix of a complex medium [14]. In such works, complex modulation is usually achieved by encoding the optical phase into the spatial displacement of binary fringes displayed on the DMD, followed by filtering high spatial frequencies in the Fourier plane [15]. Such a configuration permits multi-level complex modulation, at the cost of a reduced spatial resolution. The implementation and performance of such systems are discussed in detail in a separate tutorial [16] and are not elaborated further here. For the remainder of this paper, it will be assumed that the DMD is used for complex modulation via such a method.
@@ 2 Choosing the right DMD: Diffraction effects
A significant distinction between liquid crystal modulators and DMDs lies in the geometry of the pixel surface: The pixels of the DMD are not flat, but are mirrors that can be tilted in two different positions around a rotation axis along the diagonal of the square shape of the pixels. This difference gives rise to diffraction effects that can adversely affect both modulation quality and diffraction efficiency, defined by the fraction of the incident optical power that is redistributed into a single diffraction order. The impact of these diffraction effects is highly dependent on several factors: the wavelength of illumination, the pixel pitch, and both the incident and outgoing angles. Therefore, in conjunction with selecting an appropriate anti-reflection coating, it is crucial to ensure that the pixel pitch is compatible with the specific configuration being used. Texas Instruments offers chips with a variety of pixel pitches $d$, ranging approximately from $5$ to $\sim 25$ µm [18].
@@ 2.1 A 1D model
To achieve a qualitative understanding of this issue, we consider a 1D array of pixels as illustrated in Fig. 2. Initially, let’s assume that all pixels are in the same state and are illuminated by a plane wave originating from the far field. Under these conditions, the pixelated modulator essentially functions as a grating, with a period $d$ that is equivalent to the pixel pitch. It is important to underscore that these modulators possess a hardware-limited filling fraction, typically around 90%. This translates to an effective active pixel size of $d^{\prime}<d$.
Here we compare two types of gratings: Flat gratings, corresponding to LC-SLMs (Liquid Crystal Spatial Light Modulators), in which all pixels are in the same state; and blazed gratings, corresponding to DMDs, in which all pixels are tilted by the same angle. A grating gives rise to multiple diffraction orders, whose amplitudes are modulated by a global envelope. Importantly, these two effects can be decoupled. The diffraction order angles $\theta_{p}$ are determined by the grating periodicity, while the envelope, and in particular the position of its maximum $\theta_{\mathrm{\max}}^{envelope}$, is governed by the structure of a single pixel, which constitutes the unit cell of the grating.
2.1.1 Diffraction orders
In general, a grating gives rise to various diffraction orders with different intensities and angles $\theta_{p}$, as dictated by the grating equation:
$\text{sin}(\theta_{p})+\text{sin}(\alpha)=p\lambda/d=p\,\text{sin}(\theta_{D}),$ (1)
where $\lambda$ is the wavelength of the light, $\alpha$ is the incident angle, $\theta_{D}=\arcsin\left(\lambda/d\right)$ is the angle of the first diffraction order, and $p$ is an integer value denoting the orders of diffraction. The intensity of the individual diffraction orders is influenced by the response of a single pixel, constituting the unit cell of the grating, and that is governed by $d$, $d^{\prime}$, and its rotation angle in the case of a DMD. While this regime is not always relevant, we can adopt the small-angle approximation to obtain a qualitative picture of the effect of the incident angle $\alpha$ on the diffraction orders. In this case, we have $\theta_{p}\approx p\lambda/d-\alpha$. Thus, changing the incident angle effectively rotates all diffraction orders by an angle $\alpha$.
@@ 2.1.1 Diffraction orders
2.1.2 Position of the maximum of the envelope
The envelope of the diffracted field, which sets the amplitude of each diffraction order, is governed by the response of each unit cell, i.e. each pixel. In a flat grating excited at normal incidence, the envelope is peaked around the normal to the surface, i.e. $\theta_{max}^{envelope}=0$. In a blazed grating, the tilt angle $\theta_{B}$ introduces a linear phase ramp in the response of each individual pixel. A flat grating can be seen as a particular case with $\theta_{B}=0$. The incidence angle $\alpha$ also adds a global phase slope. When the filling fraction approaches 100%, i.e. when $d\approx d^{\prime}$, the cumulative effect leads to a shift of the envelope by an angle $2\theta_{B}-\alpha$. This results in a maximum of the envelope at
$\theta_{max}^{envelope}=2\theta_{B}-\alpha$ (2)
which reduces to $\theta_{max}^{envelope}=-\alpha$ in the case of a flat grating. We detail the calculation of this effect in Appendix A.
2.1.3 Blazed grating condition
In the case of a flat grating ($\theta_{B}=0$), we see from Eqs. 1 and 2 that the position of the first order is always aligned with the maximum of the envelope at the angle $\alpha$, when the filling fraction is close to 1. In the general case $\theta_{B}\neq 0$, $\theta_{max}^{envelope}$ does not, in general, coincide with any diffraction order [11, 13]. This leads to a redistribution of the energy of the incoming beam into multiple diffraction orders, resulting in a reduced diffraction efficiency as well as cross-talk between different pixel states, as studied in section 2.4.
The maximal diffraction efficiency is obtained when the maximum of the envelope matches one order of diffraction. This state is achieved when the conditions of the blazed grating equation are fulfilled [19], i.e. when there exist an integer $p$ that satisfies:
@@ 2.1.3 Blazed grating condition
In particular, in the Littrow configuration, i.e. for $\alpha=\theta_{B}$, we have $\theta_{max}^{envelope}=\alpha$; the diffraction and incidence angles are identical. The optimal condition is then satisfied for $\sin(\theta_{B})={p\lambda}/{(2d)}$ $p\sin(\theta_{D})/2$.
2.1.4 Example
Figure 3: 1D grating example. All angles are defined with respect to the normal of the surface of the the grating, denoted by $\vec{n}_{G}$, and are taken as positive for clockwise rotations and negative for counter-clockwise rotations. The blazing angle $\theta_{B}$ is the angle between the normal to the micro-mirror $\vec{n}_{M}$ and $\vec{n}_{G}$. The angles of the diffraction orders $\theta_{p}$ are independent of $\theta_{B}$, while the maximum of the envelope is at $\theta_{max}^{envelope}=2\theta_{B}-\alpha$.
We represent in Fig. 3 an example of a geometry for a 1D blazed grating and in Fig. 4 its angular response as well as the one corresponding to a flat grating with the same parameters. We take a 1D filling fraction of 95% (correponding to a 2D filling fraction of $\approx 90$%). For a flat grating, the zero-th order contains most of the intensity, the other orders being negligible in comparison. For the blazed grating example shown, we are in a situation close to the worst case scenario: Two diffraction orders have a significant and comparable intensity, and other orders also have non-negligible contributions. In the optimal scenario, where the peak of the envelope corresponds to a diffraction order, it results in a single diffraction order carrying the majority of the energy. This state is achieved when the conditions of the blazed grating equation are fulfilled [19]:
Figure 4: Flat grating vs blazed grating. Far field diffraction patterns for a 1D flat grating (left) and a 1D blazed grating (right) for an input angle of $\alpha=-20^{\circ}$, a filling fraction of 95% (corresponding to a 2D filling fraction of $\approx 90$%), a pixel tilt angle of $\theta_{B}=5^{\circ}$, and a wavelength to pixel pitch ratio $\lambda/d=0.05$. Vertical lines represent the angles of the diffraction orders, peaked at $\theta_{p}$ which satisfy $\sin(\theta_{p})+\sin(\alpha)=p\lambda/d$, and the black dashed curve represents the amplitude of the field which peaks at $\theta_{max}^{envelope}=2\theta_{B}-\alpha$.
@@ 2.2 The 2D case
2.2.1 Blazed grating condition
A DMD consists of a 2D array of micro-mirrors. In most DMDs, the rotation axis of the mirrors is aligned with the pixel diagonals. For convenience in the alignment and manipulation of the optical setup, it is preferable to work with incident and outgoing beams whose optical axes lie in the horizontal plane. A straightforward and common solution is to rotate the chip by $45^{\circ}$ with respect to the horizontal plane, which makes the pixels’ rotation axis vertical. This configuration is depicted in Fig. 5. We show in Appendix B. that it leads to a new blazed grating equation that reads:
Figure 5: 2D grating geometry. (a) Schematic representation of the geometry of the DMD. The incident angle $\alpha$ and the reflection angle $\theta_{max}^{envelope}$ (defined by the peak of the diffraction envelope) are situated within the horizontally plane, illustrated in yellow. (b) Photograph of the DMD chip oriented so that the rotation axis of the pixels is aligned vertically.
@@ 2.2.1 Blazed grating condition
2.2.2 Blazed number
We can quantify how close we are to the ideal case, i.e., when satisfying the blazed equation, by defining a blazed number $\mu$ as introduced in [20]:
@@ 2.2.2 Blazed number
where $\mod{2}$ represents the modulo 2 operation. $\mu$ is maximal and equals $1$ when the blazing equation is satisfied, i.e. when one order of diffraction contains most of the energy, corresponding to one diffraction order $p$ being close to the maximum of the envelope ($\theta_{p}\approx\theta^{\mathrm{envelope}}_{\max}$). $\mu$ is minimal in the worst-case scenario, i.e. when four diffraction orders have a significant and equal intensity. Note that, since the maximum of the envelope is no longer aligned with the zeroth order, corresponding to the specular reflection on the micro-mirrors, the value of the order $p$ that aligns with the envelope peak when $\mu\approx 1$ can take any integer value, and depends on the DMD parameters and on the angle of incidence $\alpha$.
2.2.3 Simulations
@@
To demonstrate the effect, we first conduct a simulation of a DMD using Python (refer to tutorial and code in [20]) with two pixel pitches of d=7.6d=7.6µm and d=10.8d=10.8µm, under a coherent excitation at $\lambda=633$nm. Fig. 6 shows the estimated blazed number $\mu$ as a function of the angle of incidence $\alpha$ in the horizontal plane, along with the far field diffraction pattern for two distinct incident angles. Fig. 7 shows the blazed number as a function of both the incident angle and the wavelength for pixel pitches of d=7.6d=7.6µm and d=10.8d=10.8µm. It should be noted that the efficiency of diffraction, i.e. the fraction of the incident optical power that is diffracted into a given diffraction order relative to the incident power, can be altered by adjusting the angle of incidence $\alpha$. However, its impact is relatively confined within an acceptable angular range that aligns with experimental limitations (i.e. for angles far from $\pm 90^{\circ}$). Far-field patterns are centered around the maximum of the envelope for an output angle $2\theta_{B}-\alpha$ (marked by a red cross). We see that for small positive values of $\alpha$, the pixel pitch of d=10.8d=10.8µm leads to a blazed number $\mu$ close to $1$. It corresponds in the far field to having one bright order of diffraction close to the maximum of the envelope.
@@
Figure 6: Blazed number and far-field diffraction patterns. Blazing number $\mu$ (Eq.(5)) as a function of the incident angle $\alpha$ (top) for a pixel pitch of d=7.6d=7.6µm (a.) and d=10.8d=10.8µm (b.). Corresponding far-field diffraction patterns (bottom) for two incident angles $\alpha=-15^{\circ}$ and $\alpha=0^{\circ}$. The red cross indicates the maximum of the envelope $\theta^{envelope}_{\text{max}}=2\theta_{B}-\alpha$ with $\theta_{B}=-12^{\circ}$.
@@
Figure 7: Blazed number as a function of the wavelength and the incident angle. Color map representation of the blazed number for a pixel pitch of d=7.6d=7.6µm (a.) and d=10.8d=10.8µm (b.) for $\theta_{B}=-12^{\circ}$.
@@
To illustrate the link between the blaze number and the efficiency of the diffraction setup, we compute the ratio of the energy in the brightest diffraction spot to the total diffracted energy for different values of the wavelength, with fixed values for the pixel pitch (7.6 µm) and incident angle (normal incidence). We simulate the DMD with no gap between the pixels (i.e., a filling fraction of 100%) and with 10 pixels in each direction, where all the mirrors of the pixels are in the same state. The results are shown in Fig. 8. We observe that the maxima (resp. minima) of the diffraction efficiency correspond to the maxima (resp. minima) of the blaze number. Note that the actual values of the minimum and maximum diffraction efficiency depend on the parameters of the DMD, such as pixel pitch, filling fraction, resolution, etc. In a practical situation, pixels are not all in the same state: the configuration is a mix of on and off states corresponding to angles $\pm\theta_{B}$. The diffraction efficiency then also depends on the modulation scheme and on the specific parameters used for the modulation. This topic is addressed in detail in a separate tutorial focusing on complex modulation using binary amplitude modulators [16].
@@
Figure 8: Blazed number vs diffraction efficiency. Plot of the blazed number and the diffraction efficiency for a normal incidence and a pixel pitch of 7.6 µm as a function of the wavelength. The diffraction efficiency is estimated by comparing the energy in a diffraction limited spot at the location of the brightest diffraction order to the total diffracted energy.